# “We Demand Justice!”: Towards Social Context Grounding of Political Texts

### v2

@@ 6.1 Ablation Analysis on Vague Text Task
We report ablation studies in table 5 on the Vague Text task test set. We consider $55$ splits: (1) Unseen Party: ⟨party, event⟩not in the train set but ⟨opposing-party, event⟩is present, (2) Unseen Event: ⟨party ⟩not in train set, (3) Flip Event: negative samples with corresponding ‘event flipped-party/vague tweet matched’ positive samples in train set and analogous (4) Flip Party and (5) Flip Tweet splits. We report the representatives from each category. We observe that all models obtain weaker performance on unseen splits, as expected, unseen events being the hardest. Contextualized models achieve higher margins. DCF gains 7.6(13.2%) and DCF embeddings attain 8.12(20.42%) macro-F1 improvement over BERT-base+wiki compared to respective margins of 8.86% and 11.42% on the full test set. In the flip splits with only negative examples, accuracy gain over random baseline for all splits is seen. This indicates that models learn to jointly condition on context information rather than learn spurious correlations over particular aspects of the context. Specifically, flip-tweet split results indicate that models don’t just learn party-explanation mapping.
@@ 6.4 Target Entity Visualization
In table 4, we study an event, Kavanaugh Supreme Court Nomination, using entity target sentiment annotations. We identify entities that are discussed by both parties. We further separate them into divisive and agreed-upon entities based on expressed sentiments. We also show partisan discussed entities for the event. This analysis paints a very accurate picture of the discussed event. We observe that the main entities of Trump, Dr. Ford, Kavanaugh, Sen. McConnell, and other accusers/survivors emerge as divisive entities. Entities such as Susan Collins and Anita Hill who were vocal mouthpieces of the respective party stances but didn’t directly participate in the event emerge as partisan entities. Supreme Court, FBI, and other entities occur in the discourse but only as neutral entities. This analysis shows the usefulness of the Target Entity and Sentiment identification in accurately summarizing events. Successful models on these tasks could greatly help in understanding political events better.
@@ 6.5 DCF Context Understanding
We look into examples that are incorrectly predicted using Wikipedia pages but correctly predicted when using the DCF model. We report some of the examples in table 6 in the appendix. In examples $11$ & $22$ of Target Entity-Sentiment task, we see that when the entity is not explicitly mentioned in the tweet, the Wiki-Context model fails to identify them as the target entities. We posit that while the Wikipedia page of each relevant event will contain these names, explicit modeling of entities in DCF model allows these examples to be correctly classified. Examples 13131-31 - 3 of Vague Text Disambiguation task show that when no clear terms indicate the sentiment towards a view, the Wiki-Context model fails to disambiguate the tweet text. Explicit modeling of politician nodes in the DCF model might allow these texts to be correctly disambiguated.
@@ 7 Conclusion and Future Work
In this paper, we motivate, define, and operationalize “Social Context Grounding” for political discourse. We build two datasets that are useful to evaluate social context grounding in NLP models. We experiment with many types of context modeling and benchmark SOTA models. We show that explicit modeling of social context outperforms naive concatenation of context and LLMs such as GPT-3. Future work includes building datasets for other components of Social Context Grounding and models that account for context in various forms.
@@
Our work only addresses English language text in US political domain. We also build upon large language models and large PLMs which are trained upon huge amounts of uncurated data. Although we employed human validation at each stage, biases could creep into the datasets. We also don’t account for the completeness of out datasets as it is a pioneering work on a new problem. Social context is vast and could have myriad of components. We only take a step in the direction of social context grounding in this work. The performance on these datasets might not indicate full social context understanding but they should help in sparking research in the direction of models that explicitly model such context. Although we tuned our prompts a lot, it is possible that better prompts and evolving models might produce better results on the LLM baselines. Our qualitative analysis is predicated on a handful of examples. They are attempts to interpret the results of large neural models and hence don’t carry as much confidence as our empirical observations.
@@
Appendix A Reproducibility
We use the HuggingFace Transformers (Wolf et al., 2020) library for PLMs. We use GPT-NeoX implementation by ElutherAI (Black et al., 2022) and GPT-3 (Brown et al., 2020b) via OpenAI API for our LLM baselines. We run $100100$ epochs for all experiments. We use $1010$ NVIDIA GeForce 1080i GPUs for our experiments. We use the train, development, and test splits detailed in section 3 for our experiments. We use the development macro-F1 for early stopping. We run all our experiments using random seeds to ensure reproducibility. We experiment with a random seed value set to {13131313}. We report the results of the $33$ fold cross-validation. We report only the mean on all the cross-fold validation results for clarity. All our code, datasets, and result logs will be released publicly upon acceptance. We experiment with $33$, $55$, and $1010$ fold cross-validation. As the results on the development data are almost identical, we report the results of $33$ fold cross-validation in all our experiments.
Appendix B Error Analysis
@@
Appendix C Annotation Interfaces

### v3

@@ 6.1 Ablation Analysis on Vague Text Task
We report ablation studies in table 5 on the Vague Text task test set. We consider $55$ splits: (1) Unseen Party: ⟨party, event⟩not in the train set but ⟨opposing-party, event⟩is present, (2) Unseen Event: ⟨party ⟩not in train set, (3) Flip Event: negative samples with corresponding ‘event flipped-party/vague tweet matched’ positive samples in train set and analogous (4) Flip Party and (5) Flip Tweet splits. We observe the best model in each category. They obtain weaker performance on unseen splits, as expected, unseen events being the hardest. Contextualized models achieve higher margins. DCF gains 7.6(13.2%) and DCF embeddings attain 8.12(20.42%) macro-F1 improvement over BERT-base+wiki compared to respective margins of 8.86% and 11.42% on the full test set. In the flip splits with only negative examples, accuracy gain over random baseline for all splits is seen. This indicates that models learn to jointly condition on context information rather than learn spurious correlations over particular aspects of the context. Specifically, flip-tweet split results indicate that models don’t just learn party-explanation mapping.
@@ 6.4 Target Entity Visualization
The main goal of this analysis is to demonstrate the usefulness and inspire modeling research in the direction of entity-sentiment-centric view of political events. table 4 visualizes one component of how partisan discourse is structured in these events. We study Kavanaugh Supreme Court Nomination. We identify discussed entities and separate them into divisive and agreed-upon entities. This analysis paints an accurate picture of the discussed event. We observe that the main entities of Trump, Dr. Ford, Kavanaugh, Sen. McConnell, and other accusers/survivors emerge as divisive entities. Entities such as Susan Collins and Anita Hill who were vocal mouthpieces of the respective party stances but didn’t directly participate in the event emerge as partisan entities. Supreme Court, FBI, and other entities occur but only as neutral entities.
@@ 6.5 DCF Context Understanding
We look into examples that are incorrectly predicted using Wikipedia pages but correctly predicted by the DCF model in the appendix (table 6). In examples $11$ & $22$ of Target Entity-Sentiment task, when the entity is not explicitly mentioned in the tweet, the Wiki-Context model fails to identify them as the targets. We posit that while the Wikipedia page of each relevant event will contain these names, explicit modeling of entities in the DCF model allows correct classification. Examples 13131-31 - 3 of Vague Text Disambiguation task show that when no clear terms indicate the sentiment towards a view, the Wiki-Context model fails to disambiguate the tweet text. Explicit modeling of politician nodes seems to help the DCF model.
@@ 7 Conclusion and Future Work
In this paper, we motivate, define, and operationalize ‘Social Context Grounding’ for political text. We build two novel datasets to evaluate social context grounding in NLP models that ‘are easy for humans’ when the relevant social context is provided. We experiment with many types of contextual models. We show that explicit modeling of social context outperforms other models while lacking behind humans.
Acknowledgements
We thank Shamik Roy, Nikhil Mehta, and the anonymous reviewers for their vital feedback. The project was funded by NSF CAREER award IIS-2048001 and the DARPA CCU Program. The contents are those of the author(s) and do not necessarily represent the official views of, nor an endorsement by, DARPA, or the US Government.
@@ Acknowledgements
Our work only addresses English language text in US political domain. We also build upon large language models and large PLMs which are trained upon huge amounts of uncurated data. Although we employed human validation at each stage, biases could creep into the datasets. We also don’t account for the completeness of our datasets as it is a pioneering work on a new problem. Social context is vast and could have a myriad of components. We only take a step in the direction of social context grounding in this work. The performance on these datasets might not indicate full social context understanding but they should help in sparking research in the direction of models that explicitly model such context. Although we tuned our prompts a lot, better prompts and evolving models might produce better results on the LLM baselines. Our qualitative analysis is predicated on a handful of examples. They are attempts to interpret the results of large neural models and hence don’t carry as much confidence as our empirical observations. We believe the insights from our findings will encourage more research in this area. For example, the development of discourse contextualized models that aim to model human-style understanding of background knowledge, emotional intelligence, and societal context understanding is a natural next step of our research.
@@
Appendix A GPT Prompts
Prompts for Target-Entity Task:
Event: <event>
Event background: <background-description>
Tweet: <tweet-text>
Author: <author-name>
Author Party: <party-affiliation>
Author background: <first two sentences of author-wiki-page>
Target Entity: <entity-name>
Entity background: <first two sentences of entity-wiki-page>
Task: Identify if the given entity is a target of the tweet. A target entity is defined as an entity that would be present in the full unambiguous explanation of the tweet.
Is the given entity a target entity of the tweet? Answer yes or no.
Prompts for Target-Sentiment Task:
Event: <event>
Event background: <background-description>
Tweet: <tweet-text>
Author: <author-name>
Author Party: <party-affiliation>
Author background: <first two sentences of author-wiki-page>
Target Entity: <entity-name>
Entity background: <first two sentences of entity-wiki-page>
Task: Identify the sentiment of the tweet towards the given target entity. Consider that the tweet is ambiguous and the entity might be implied without being explicitly mentioned.
What is the sentiment of the tweet towards the target entity? Answer with positive, negative, or neutral.
Prompts for Vague Text Task:
Event: <event>
Event background: <background-description>
Vague message: <vague-text>
Author Party: <party-affiliation>
Author background: <first two sentences of party-wiki-page>
Task: Given the event, vague message, and party affiliation of the author, explain unambiguously the intended meaning of the vague message.
Generate an unambiguous explanation for the vague message given the party affiliation of the author and the event in context.
Appendix B Reproducibility
We use the HuggingFace Transformers (Wolf et al., 2020) library for PLMs. We use GPT-NeoX implementation by ElutherAI (Black et al., 2022) and GPT-3 (Brown et al., 2020b) via OpenAI API for our LLM baselines. We run $100100$ epochs for all experiments. We use $1010$ NVIDIA GeForce 1080i GPUs for our experiments. We use the train, development, and test splits detailed in section 3 for our experiments. We use the development macro-F1 for early stopping. We run all our experiments using random seeds to ensure reproducibility. We experiment with a random seed value set to {13131313}. We set CUBLAS environment variables for reproducibility. All our code, datasets, and result logs are released publicly.
Appendix C Error Analysis
@@
Appendix D Annotation Interfaces